Alex travels twice as far as john, but only one third as fast. John starts at point x at noon on Monday and travels to point z, 300 away, b y 9:00 pm the same day. If alex starts 3 hour later than john, when will alex reach point z.. ( assuming no stops were made

Answer by addingup(3677)

Distance= rate x time (rate is another word for speed)
For John, we know the distance (300mi) and the time (9hrs) so we need to find the rate:
We divide both sides of the equation by time:
Distance/time= rate
300/9= 33.33 miles per hour is John's speed.
----------------------
Alex travels twice as far at 1/3 the speed, so Alex goes 600 miles at 11.11mph:
Distance= rate*time Divide both sides by rate
distance/rate= time
600/11.11= 54hours; 54/24= 2 days and 6 hours
Alex starts at 12+3 at 3pm and finishes 2 days later at 9pm